### v2

@@ B.2.1 Longest Path Problem
$\begin{aligned} \min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad&\quad\textnormal{loss}(\boldsymbol{p}_{\mathcal{K}},\boldsymbol{x}_{\mathcal{K}})\\ \text{s.t.}\quad\ -&\delta^{k,l}_{i,j}\ \leq\ \lambda_{j}^{k}-\lambda_{i}^{k}-\lambda_{j}^{l}+\lambda_{i}^{l}\ \leq\ 1-(1+\epsilon)\delta^{k,l}_{i,j},\quad\ \ \forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\delta^{k,l}_{i,j}-1\ \leq\ x_{ij}^{l}-x_{ij}^{k}\ \leq\ 1-\delta^{l,k}_{i,j},\quad\qquad\quad\forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ -&(\delta^{k,l}_{i,j}+\delta^{l,k}_{i,j})\ \leq\ x_{ij}^{l}-x_{ij}^{k}\ \leq\ \delta^{k,l}_{i,j}+\delta^{l,k}_{i,j},\ \quad\forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\sum_{j:(i,j)\in\boldsymbol{E}^{k}}x_{ij}-\sum_{j:(j,i)\in\boldsymbol{E}^{k}}x_{ji}=\begin{cases}1,\quad&i=s,\\ -1,\quad&i=t,\\ 0,\quad&i\in\boldsymbol{V}^{k},\end{cases}\forall k\in\mathcal{K},\\ &\delta^{k,l}_{i,j}\,\in\,\{0,1\},\ \forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l}.\end{aligned}$ (16)
B.2.2 Shortest Path Problem
@@
S-RAM characterization for Shortest Path Problems. Under Assumption 1, a data collection $\boldsymbol{p}^{\mathcal{K}}$ is representable by an S-RAM if and only if there exists $(\lambda_{i}^{k}\in\mathbb{R}:i\in s\cup t\cup\boldsymbol{V}^{k},k\in\mathcal{K})$, such that for any two distinct graph $k,l\in\mathcal{K}$ containing a common edge $(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},$
$\begin{aligned} &\lambda_{j}^{k}-\lambda_{i}^{k}<\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad p_{ij}^{k}<p_{ij}^{l},\\ &\lambda_{j}^{k}-\lambda_{i}^{k}=\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1.\end{aligned}$ (17)
@@
$\begin{aligned} \max_{\epsilon,\boldsymbol{\lambda}}\quad&\epsilon\\ \text{s.t.}\quad&\lambda_{j}^{k}-\lambda_{i}^{k}+\epsilon\leq\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad p_{ij}^{k}<p_{ij}^{l},\quad\forall k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\lambda_{j}^{k}-\lambda_{i}^{k}=\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1,\quad\forall k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l}.\end{aligned}$ (18)
@@
$\min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad$ $\quad f(\boldsymbol{x})$ (19)
@@
Likewise, the best-case estimate for the objective value $\bar{f}_{\text{new}}$ equals the optimal value obtained by maximizing over the constraints in the above mixed integer linear program. In our experiment, we take $f(\boldsymbol{x})=x_{\hat{e}}$ where $\hat{e}$ is the pre-specified edge.
@@
$\begin{aligned} \min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad&\quad\textnormal{loss}(\boldsymbol{p}_{\mathcal{K}},\boldsymbol{x}_{\mathcal{K}})\\ \text{s.t.}\quad\ -&\delta^{k,l}_{i,j}\ \leq\ \lambda_{j}^{k}-\lambda_{i}^{k}-\lambda_{j}^{l}+\lambda_{i}^{l}\ \leq\ 1-(1+\epsilon)\delta^{k,l}_{i,j},\quad\ \ \forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\delta^{k,l}_{i,j}-1\ \leq\ x_{ij}^{l}-x_{ij}^{k}\ \leq\ 1-\delta^{l,k}_{i,j},\quad\qquad\quad\,\forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ -&(\delta^{k,l}_{i,j}+\delta^{l,k}_{i,j})\ \leq\ x_{ij}^{l}-x_{ij}^{k}\ \leq\ \delta^{k,l}_{i,j}+\delta^{l,k}_{i,j},\ \quad\forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\sum_{j:(i,j)\in\boldsymbol{E}^{k}}x_{ij}-\sum_{j:(j,i)\in\boldsymbol{E}^{k}}x_{ji}=\begin{cases}1,\quad&i=s,\\ -1,\quad&i=t,\\ 0,\quad&i\in\boldsymbol{V}^{k},\end{cases}\forall k\in\mathcal{K},\\ &\delta^{k,l}_{i,j}\,\in\,\{0,1\},\ \forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l}.\end{aligned}$ (20)
B.3 Assignment Problem
@@
S-RAM characterization for Assignment Problems. Under Assumption 1, an observed flow collection $\boldsymbol{p}^{\mathcal{K}}$ is representable by an S-RAM if and only if there exists $(\lambda_{i,L}^{k},\lambda_{j,R}^{k}\in\mathbb{R}:i\in\boldsymbol{V}^{k}_{L},j\in\boldsymbol{V}^{k}_{R},k\in\mathcal{K})$, such that for any two distinct graph $k,l\in\mathcal{K}$ containing a common edge $(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},$
$\begin{aligned} &\lambda_{i,L}^{k}+\lambda_{j,R}^{k}>\lambda_{i,L}^{l}+\lambda_{j,R}^{l}\quad\text{if}\quad p_{ij}^{k}>p_{ij}^{l},\\ &\lambda_{i,L}^{k}+\lambda_{j,R}^{k}=\lambda_{i,L}^{l}+\lambda_{j,R}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1.\end{aligned}$ (21)
@@
$\begin{aligned} \max_{\epsilon,\boldsymbol{\lambda}}\quad&\epsilon\\ \text{s.t.}\quad&\lambda_{i,L}^{k}+\lambda_{j,R}^{k}\geq\lambda_{i,L}^{l}+\lambda_{j,R}^{l}+\epsilon\quad\text{if}\quad p_{ij}^{k}>p_{ij}^{l},\ \forall k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\lambda_{i,L}^{k}+\lambda_{j,R}^{k}=\lambda_{i,L}^{l}+\lambda_{j,R}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1,\ \forall k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l}.\end{aligned}$ (22)
@@
$\min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad$ $\quad f(\boldsymbol{x})$ (23)
@@
$\sum_{j\in\boldsymbol{V}_{R}^{k}}x_{ij}=1,\quad\forall k\in\mathcal{K},\forall i\in\boldsymbol{V}_{L}^{k},$ (24)

### v3

@@ B.1. Data Generation
(19) $\begin{aligned} \min_{\boldsymbol{x}}&\sum_{(i,j)\in\boldsymbol{E}^{k}}(\nu_{ij}+a_{ij})x_{ij}+\sum_{(i,j)\in\boldsymbol{E}^{k}}\frac{b_{ij}-a_{ij}}{2}x_{ij}^{2}\\ \text{s.t.}\;&\sum_{j:(i,j)\in\boldsymbol{E}^{k}}x_{ij}-\sum_{j:(j,i)\in\boldsymbol{E}^{k}}x_{ji}=\begin{cases}1,\quad&i=s,\\ -1,\quad&i=t,\\ 0,\quad&i\in\boldsymbol{V}^{k},\end{cases}\\ &x_{ij}\geq 0\quad\forall\ (i,j)\in\boldsymbol{E}^{k}.\end{aligned}$
These choice probabilities are used as the observed data $\boldsymbol{p}^{\mathcal{K}}$ to train the nonparametric s-RAM model. For prediction, we generate a new graph $\boldsymbol{G}^{\text{new}}$ and compute the probabilities via 19 with $\boldsymbol{G}^{\text{new}}$.
Assignment Problem. Recall that for the assignment problem, we assume the noise marginals are nonidentical uniform distributions. Using the parameter notations in the shortest path problem above, from (Natarajan et al., 2009), we have that, for a graph $\boldsymbol{G}^{k}$ with $k\in\mathcal{K}$, the convex program with nonidentical uniform marginal distributions for the assignment problem is as follows:
(20) $\begin{aligned} \max_{\boldsymbol{x}}&\sum_{i\in\boldsymbol{V}_{L}^{k}}\sum_{j\in\boldsymbol{V}_{R}^{k}}(\nu_{ij}+b_{ij})x_{ij}-\sum_{i\in\boldsymbol{V}_{L}^{k}}\sum_{j\in\boldsymbol{V}_{R}^{k}}\frac{b_{ij}-a_{ij}}{2}x_{ij}^{2}\\ \text{s.t.}\;&\sum_{j\in\boldsymbol{V}_{R}^{k}}x_{ij}=1,\quad\forall i\in\boldsymbol{V}_{L}^{k},\\ &\sum_{i\in\boldsymbol{V}_{L}^{k}}x_{ij}=1,\quad\forall j\in\boldsymbol{V}_{R}^{k},\\ &x_{ij}\geq 0\quad\forall\ i\in\boldsymbol{V}_{L}^{k},j\in\boldsymbol{V}_{R}^{k}.\end{aligned}$
These are used as observed data $\boldsymbol{p}^{\mathcal{K}}$ for training the nonparametric s-RAM model. For prediction, we generate a new graph $\boldsymbol{G}^{\text{new}}$ and compute the total expected system welfare via 20 with $\boldsymbol{G}^{\text{new}}$, which is the optimal objective value under the new graph.
Constrained Subset Selection Problem. Recall that for the constrained subset selection problem, we assume the noise marginals are nonidentical exponential distributions. Given a collection historical subsets $\mathcal{I}^{(1)},\mathcal{I}^{(2)},\cdots,\mathcal{I}^{(K)}$, for a subset $\mathcal{I}^{(k)}$ with $k\in\mathcal{K}$, we solve a convex program defined by the corresponding parametric MDM model to obtain the choice probabilities. Let $\boldsymbol{\nu}$ denote the deterministic weight matrix. Let $\lambda_{i}$ be the rate parameters for each item $i\in\mathcal{N}$. From (Natarajan et al., 2009), we have that, for a subset $\mathcal{I}^{(k)}$, the convex program with nonidentical exponential marginal distributions for the constrained subset selection problem is as follows:
(21) $\begin{aligned} \max_{\boldsymbol{x}}&\sum_{i\in\mathcal{I}^{(k)}}(\nu_{i}+\frac{1}{\lambda_{i}})x_{i}-\sum_{i\in\mathcal{I}^{(k)}}\frac{1}{\lambda_{i}}x_{i}\ln{x_{i}}\\ \text{s.t.}\;&\sum_{i\in\mathcal{I}^{(k)}}x_{i}\leq C,\\ &0\leq x_{ij}\leq 1\quad\forall\ i\in\mathcal{I}^{(k)}.\end{aligned}$
These are used as observed data $\boldsymbol{p}^{\mathcal{K}}$ for training the nonparametric s-RAM model. For prediction, we generate a new graph $\boldsymbol{G}^{\text{new}}$ and compute the marginal allocation weights via 21 with $\boldsymbol{G}^{\text{new}}$, which is the probability of the chosen item.
B.2. Representability, Estimation, Prediction Formulations
B.2.1. Shortest Path Problem
@@
s-RAM characterization for Shortest Path Problems. Under Assumptions 1 and 2, a data collection $\boldsymbol{p}^{\mathcal{K}}$ is representable by an s-RAM if and only if there exists $(\lambda_{i}^{k}\in\mathbb{R}:i\in s\cup t\cup\boldsymbol{V}^{k},k\in\mathcal{K})$, such that for any two distinct graph $k,l\in\mathcal{K}$ containing a common edge $(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},$
(22) $\begin{aligned} &\lambda_{j}^{k}-\lambda_{i}^{k}<\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad p_{ij}^{k}<p_{ij}^{l},\\ &\lambda_{j}^{k}-\lambda_{i}^{k}=\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1.\end{aligned}$
@@
(23) $\begin{aligned} \max_{\epsilon,\boldsymbol{\lambda}}\quad&\epsilon\\ \text{s.t.}\quad&\lambda_{j}^{k}-\lambda_{i}^{k}+\epsilon\leq\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad p_{ij}^{k}<p_{ij}^{l},\quad\forall k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\lambda_{j}^{k}-\lambda_{i}^{k}=\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1,\quad\forall k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l}.\end{aligned}$
@@
(24) $\min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad$ $\quad f(\boldsymbol{x})$
@@
Likewise, the best-case estimate for the objective value $\bar{f}_{\text{new}}$ equals the optimal value obtained by maximizing over the constraints in the above mixed integer program. In our experiment, we take $f(\boldsymbol{x})=x_{\hat{e}}$ where $\hat{e}$ is the pre-specified edge.
@@
(25) $\begin{aligned} \min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad&\quad\textnormal{loss}(\boldsymbol{p}_{\mathcal{K}},\boldsymbol{x}_{\mathcal{K}})\\ \text{s.t.}\quad\ -&\delta^{k,l}_{i,j}\ \leq\ \lambda_{j}^{k}-\lambda_{i}^{k}-\lambda_{j}^{l}+\lambda_{i}^{l}\ \leq\ 1-(1+\epsilon)\delta^{k,l}_{i,j},\quad\ \ \forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\delta^{k,l}_{i,j}-1\ \leq\ x_{ij}^{l}-x_{ij}^{k}\ \leq\ 1-\delta^{l,k}_{i,j},\quad\qquad\quad\,\forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ -&(\delta^{k,l}_{i,j}+\delta^{l,k}_{i,j})\ \leq\ x_{ij}^{l}-x_{ij}^{k}\ \leq\ \delta^{k,l}_{i,j}+\delta^{l,k}_{i,j},\ \quad\forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\sum_{j:(i,j)\in\boldsymbol{E}^{k}}x_{ij}-\sum_{j:(j,i)\in\boldsymbol{E}^{k}}x_{ji}=\begin{cases}1,\quad&i=s,\\ -1,\quad&i=t,\\ 0,\quad&i\in\boldsymbol{V}^{k},\end{cases}\forall k\in\mathcal{K},\\ &x_{ij}\geq 0,\ \forall\,k\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k},\\ &\delta^{k,l}_{i,j}\,\in\,\{0,1\},\ \forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l}.\end{aligned}$
B.2.2. Assignment Problem
@@
s-RAM characterization for Assignment Problems. Under Assumptions 1 and 2, an observed flow collection $\boldsymbol{p}^{\mathcal{K}}$ is representable by an s-RAM if and only if there exists $(\lambda_{i,L}^{k},\lambda_{j,R}^{k}\in\mathbb{R}:i\in\boldsymbol{V}^{k}_{L},j\in\boldsymbol{V}^{k}_{R},k\in\mathcal{K})$, such that for any two distinct graph $k,l\in\mathcal{K}$ containing a common edge $(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},$
(26) $\begin{aligned} &\lambda_{i,L}^{k}+\lambda_{j,R}^{k}>\lambda_{i,L}^{l}+\lambda_{j,R}^{l}\quad\text{if}\quad p_{ij}^{k}>p_{ij}^{l},\\ &\lambda_{i,L}^{k}+\lambda_{j,R}^{k}=\lambda_{i,L}^{l}+\lambda_{j,R}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1.\end{aligned}$
@@
(27) $\begin{aligned} \max_{\epsilon,\boldsymbol{\lambda}}\quad&\epsilon\\ \text{s.t.}\quad&\lambda_{i,L}^{k}+\lambda_{j,R}^{k}\geq\lambda_{i,L}^{l}+\lambda_{j,R}^{l}+\epsilon\quad\text{if}\quad p_{ij}^{k}>p_{ij}^{l},\ \forall k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\lambda_{i,L}^{k}+\lambda_{j,R}^{k}=\lambda_{i,L}^{l}+\lambda_{j,R}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1,\ \forall k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l}.\end{aligned}$
@@
(28) $\min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad$ $\quad f(\boldsymbol{x})$
@@
(30) $\sum_{j\in\boldsymbol{V}_{R}^{k}}x_{ij}=1,\quad\forall k\in\mathcal{K},\forall i\in\boldsymbol{V}_{L}^{k},$
@@
B.2.3. Constrained Subset Selection Problem
Suppose $\mathcal{N}$ is the complete set of available items. We focus on the cardinality-constrained case with $\mathcal{X}=\{\boldsymbol{x}:\sum_{i\in\mathcal{N}}x_{i}\leq C,\ x_{i}\in\{0,1\},\forall i\in\mathcal{N}\}$, where $C$ is a positive integer such that $2\leq C\leq|\mathcal{N}|-1$. We have marginal allocation weight data for a collection of $K$ observations with available subsets $\mathcal{I}^{(1)},\mathcal{I}^{(2)},\ldots,\mathcal{I}^{(K)}$ such that $\mathcal{I}^{(k)}\subseteq\mathcal{N}$, for $k\in\mathcal{K}$. For a set $\mathcal{I}^{(k)}$, let $p_{i}^{k}\in[0,1]$ denote the marginal allocation weight for item $i$ in $\mathcal{I}^{(k)}$. Given observed marginal allocation weight data $\boldsymbol{p}^{\mathcal{K}}=(p_{i}^{k}:i\in\mathcal{I}^{(k)},k\in\mathcal{K})$, by Theorem 3.1, the s-RAM characterization for the constrained subset selection problem is given below.
s-RAM characterization for Constrained Subset Selection Problems with Cardinality Constraint. Under Assumptions 1 and 2, a data collection $\boldsymbol{p}^{\mathcal{K}}$ is representable by an s-RAM if and only if there exists $(\lambda^{k}\in\mathbb{R}:k\in\mathcal{K})$ such that for any two observation $k,l\in\mathcal{K}$ containing a common item $i\in\mathcal{N},$
(32) $\begin{aligned} &\lambda_{k}>\lambda_{l}\quad\text{if}\quad p_{i}^{k}<p_{i}^{l},\\ &\lambda_{k}>\lambda_{l}\quad\text{if}\quad p_{i}^{k}=p_{i}^{l}\neq 0.\end{aligned}$
As a result, checking whether the given choice data $\boldsymbol{p}^{\mathcal{K}}$ satisfies the s-RAM hypothesis can be accomplished by solving a linear program with $\mathcal{O}(|\mathcal{K}|)$ continuous variables and $\mathcal{O}(n|\mathcal{K}|)$ constraints as follows, which is equivalent to checking whether the optimal objective value of (33) below is strictly positive.
(33) $\begin{aligned} \max_{\epsilon,(\lambda_{k}:k\in\mathcal{K})}\quad&\epsilon&&&&&\\ \text{s.t.}\qquad&\lambda_{k}\geq\lambda_{l}+\epsilon&&\text{if }\quad p_{i}^{k}<p_{i}^{l},&&&\forall k,l\in\mathcal{K},i\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)},\\ &\lambda_{k}=\lambda_{l}&&\text{if }\quad p_{i}^{k}=p_{i}^{l}\neq 0,&&&\forall k,l\in\mathcal{K},i\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)}.\end{aligned}$
Prediction for Constrained Subset Selection Problems. Suppose $\boldsymbol{p}^{\mathcal{K}}$ is representable by s-RAM, given a new set $\mathcal{I}^{\text{new}}$, by Proposition 4.2, the prediction formulation to compute $\underline{f}_{\text{new}}$ for the constrained subset selection problem by taking any $0<\epsilon<1/(2|\mathcal{K}|)$ is as follows:
[…]
Likewise, the best-case estimate for the objective value $\bar{f}_{\text{new}}$ equals the optimal objective value obtained by maximizing over the constraints in the above mixed integer program.
Estimation for Constrained Subset Selection Problems. Suppose that Assumptions 1 and 2 is satisfied. Then for any $0<\epsilon<1/(2|\mathcal{K}|),$ formulation 12 can be written as:
(35) $\begin{aligned} &\min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}}\quad\textnormal{loss}(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x}^{\mathcal{K}})\\ &\text{s.t.}\quad-\delta_{k,l}\leq\lambda_{k}-\lambda_{l}\leq 1-(1+\epsilon)\delta_{k,l},\quad\quad\quad\;\;\;\forall k,l\in\mathcal{K},i\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)},\\ &\quad\quad\quad\delta_{k,l}-1\leq x_{i}^{k}-x_{i}^{l}\leq 1-\delta_{l,k},\quad\quad\quad\quad\ \forall k,l\in\mathcal{K},i\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)},\\ &\quad\quad-(\delta_{k,l}+\delta_{l,k})\leq x_{i}^{k}-x_{i}^{l}\leq\delta_{k,l}+\delta_{l,k},\quad\,\forall k,l\in\mathcal{K},i\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)},\\ &\quad\quad\quad\sum_{i\in\mathcal{I}^{(k)}}x_{i}^{k}\leq C,\ \forall\,k\in\mathcal{K},\;0\leq x_{i}^{k}\leq 1,\ \forall\,k\in\mathcal{K},i\in\mathcal{I}^{(k)},\\ &\quad\quad\quad 0\leq\lambda_{k}\leq 1,\ \forall\,k\in\mathcal{K},\;\\ &\quad\quad\quad\delta_{k,l}\in\{0,1\},\ \forall\,k,l\in\mathcal{K}.\end{aligned}$